An electric hot plate has a power rating of 150w and a current draw of 10A,what is the resistance of the hot plate


1
Expert's answer
2021-02-22T10:22:19-0500

From the definition of the power, we have:


"P=I^2R,""R=\dfrac{P}{I^2}=\dfrac{150\ W}{(10\ A)^2}=1.5\ \Omega."